# Finslerian structure of black holes

### v1

@@ Abstract
Recently, there has been an increasing interest in the Finslerian interpretation of null geodesics in the exterior regions of stationary black holes, particularly through the Zermelo navigation problem and the Randers metric. In this work, we show that recent mathematical advancements in the critical and strong Zermelo navigation problems and their connections to Kropina and Lorentz-Finsler metrics enable the extension of this Finslerian framework to encompass horizons and their interior regions. The Finslerian indicatrix, a key element of this framework, serves as an effective tool for identifying frame-dragging effects and the location of horizons and ergosurfaces. We illustrate our results with explicit physical examples, focusing on spherically symmetric and Kerr black holes. Our findings provide new insights into the “river model” of black holes, offering enhanced visual representations of null geodesics on the ergosurfaces, horizons, and within their interior regions.
@@ I Introduction
where $\mathbf{x}$ is a point with coordinates $(x^{\sigma})$ in the spacetime manifold $\mathcal{M}$ and $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$. We denote the components of $d\mathbf{x}$ by $(dx^{\mu})$, following the standard physical convention in which Greek indices run over all $n+1$ spacetime coordinates, while Latin indices are limited to the $n$ spatial dimensions. In the mathematical literature, $F(\mathbf{x},d\mathbf{x})$ is typically referred to as a (Minkowski) norm rather than a metric. In this work, however, we refer to it as a metric, following Randers’ original terminology and ensuring consistency with the conventions commonly adopted in the physical literature. In the Randers metric (1), $a_{\mu\nu}$ is a standard metric tensor (Lorentzian, in the case of Spacetime Physics), and $b_{\mu}$ is a one-form (covariant vector) that encodes a directional dependence of the geometry. If $b_{\mu}=0$, the underlying geometry reduces to the standard Riemannian/Lorentzian one associated with the metric tensor $a_{\mu\nu}$. Notice that $F(\mathbf{x},d\mathbf{x})=F(\mathbf{x},-d\mathbf{x})$ for generic vectors $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$ if, and only if, $b_{\mu}=0$. In other words, a Finsler metric is symmetric if and only if it is Riemannian. Notwithstanding, generic Finsler metrics are positively homogeneous of degree one, i.e., $F(\mathbf{x},\lambda d\mathbf{x})=\lambda F(\mathbf{x},d\mathbf{x})$ for all $\lambda\geq 0$. All Finslerian-geometric objects derived from (1), such as the fundamental form, connections, covariant derivative, and curvature, are defined starting from $F^{2}(\mathbf{x},d\mathbf{x})$ analogously to the Riemannian case (see [2] and references therein for further details). Positive definite metrics of the Randers type (1) arise naturally [5] in the classical optimization problem known as the Zermelo navigation problem [6], which consists in finding minimal (non-relativistic) time trajectories in curved spaces under the influence of vector drifts (winds). Many recent advancements in Finsler geometry and its applications in Physics are closely related to the Zermelo problem and its solutions involving metrics such as those of Randers type. For a comprehensive review on the the recent mathematical literature, see [7]. This work will further explore the different types of Zermelo navigation problems and their connections to other Finslerian metrics in Spacetime Physics.
@@
From a physical standpoint, one does not expect distinguished local spatial directions. Hence, one would expect the emergence of Finslerian metrics like (1) in Spacetime Physics only when treating problems where the temporal direction is explicitly involved. In fact, this was the main motivation behind Randers’ introduction of a spacetime metric with asymmetrical properties of the type (1) in the early forties [4]. Moreover, the Principle of Equivalence imposes restrictions on possible forms of a Finsler metric from the Spacetime Physics perspective [8]. Interestingly, positive definite metrics of type (1) often appear in specific spacetime problems restricted to spatial sections, such as the description of causal structures. In particular, the propagation of light rays in stationary spacetimes can be effectively described by a Randers spatial metric. These issues were carefully and comprehensively considered for black holes in [9]. The extension for nonstationary but stably causal spacetimes admitting a globally defined time coordinate was considered in [10]. More recently, [11] has shown that the Zermelo problem can be used to generate spacetimes with specific causal properties. To date, only the propagation of light rays in the exterior region of black holes has been described in Finslerian terms in the Physics literature. Extending the Finsler structure across the event horizon is a physically relevant problem, as recognized in [11]. We highlight that recent applications of Finsler geometry in other problems of General and Special Relativity can be found, e.g., in [12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
In this paper, we demonstrate how recent mathematical progress [7, 21, 22, 23] can be explored to interpret light propagation within the interior region of black holes in Finslerian terms. The key ingredients are the Kropina and Lorentz-Finsler metrics, which appear in certain extended versions of the Zermelo navigation problem. We show, through explicit and physically relevant examples, that the Kropina metric adequately describes the causal properties of ergosurfaces in stationary black holes, while the Lorentz-Finsler metric captures, in the region where it is well defined, the causal structure of the interior (non-stationary) region of black holes. In this way, we show that it is indeed possible to have, based on recent mathematical developments, a complete Finslerian description of the causal structure of black holes, shedding new light on phenomena such as the one-directional membrane behavior of horizons [24] and the river model paradigm of black holes [25].
@@
The paper is organized as follows. In the next section, we review recent mathematical developments related to the Kropina and Lorentz-Finsler metrics and compile the results that are instrumental in describing the causal structure of black holes, which is explored in Section III. The final section is dedicated to our concluding remarks, which include brief comments on Fermat’s principle and applications to analog spacetimes. We adopt $G=c=1$ units throughout this work.
@@ II Finsler geometry and the Zermelo problem
where $a^{ij}$ is the inverse of the metric tensor $a_{ij}$. Hence, the weak drift also implies $||b||^{2}_{F}<1$, which is commonly known as the Finsler condition in the Physics literature. As it will become clear in the next section, describing the Finslerian structure of black holes requires extending the Zermelo problem to the critical $(||W||=1)$ and strong $(||W||>1)$ drift cases. Fortunately, both problems have been thoroughly investigated and, for the purposes of this work, conclusively solved in the recent mathematical literature, see [7] for a recent review on this subject.
@@
The strong drift case, on the other hand, was solved in [23], where it is shown that minimal time trajectories correspond to the geodesics of the so-called Lorentz-Finsler metrics
@@
We shall refer to the minus and plus cases, respectively, as first and second Lorentz-Finsler metrics. The strong drift condition $\lambda<0$ guarantees that both metrics in (6) are positive definite. Additionally, as in the critical case, the problem exhibits a one-directional behavior: a solution exists only if the trajectories are positively aligned with the drift, and moreover their instantaneous velocities lie within the conic region (7).
@@
In other words, the indicatrices of the Finsler metrics associated with the Zermelo navigation problem on a Riemannian manifold $(\mathcal{M},h_{ij})$ with drift $W^{i}$, regardless of the value of $||W||$, correspond to the unit balls of the Riemannian metric $h_{ij}$, with their centers shifted by the vector $W^{i}$. Randers had already recognized this type of eccentricity in his pioneering work [4]. This result will be essential for the analysis of the causal structure of black holes in Section III.
@@ II.1 Geodesic matching
The mathematical description of the causal structure of Finsler spaces is both comprehensive and sophisticated, see, e.g.[7, 26, 27, 28, 29]. For our purposes here, we need only consider the matching of geodesics across the three cases of Zermelo’s problems. Differentiability issues will not be discussed, as all functions are assumed to be sufficiently smooth. Let $W^{i}$ be a generic smooth vector field drift on a Riemannian manifold $(\mathcal{M},h_{ij})$ and assume there are regions of $\mathcal{M}$ characterized by weak ($||W||<1$) and strong $(||W||>1)$ drifts. The boundary between these regions is a co-dimension 1 submanifold $\mathcal{H}$ of $\mathcal{M}$, corresponding to the critical drift case $||W||=1$. The key observation is that geodesics can cross smoothly from the weak to the strong drift regions, provided that the half-space (5) and conic (7) constraints are satisfied in the pertinent regions. This is possible because the Randers (1), Kropina (4), and first Lorentz-Finsler (6) metrics all give rise to the same Finslerian Lagrangian action $S=\int F(\mathbf{x},d\mathbf{x})$ for geodesics. Notice that the half-space and conic constraints are essential to assure that we have positive Lagrangians $F(\mathbf{x},d\mathbf{x})$ and consequently well-defined geodesic problems, on $\mathcal{H}$ and in the strong drift region.
@@
Figure 1: Depiction of the Finslerian indicatrices for the three drift regimes of the Schwarzschild case along the $x$ axis in the equatorial plane. In all cases, the indicatrices are unit spheres, with their centers displaced in the $x$ direction according to (10). Right: The weak drift regime $r>2M$, in which null geodesics starting from a given point can propagate in all directions. Center: The critical regime $r=2M$, in which the propagation of null geodesics is restricted to the directions that are positively aligned with the drift according to the half-space restriction (5). In this case, the half-space is tangent to the critical submanifold $\mathcal{H}$ (the event horizon), implying that there is no outgoing null geodesics starting on $\mathcal{H}$. Left: The strong drift regime. Here, propagation is similarly limited to directions positively aligned with the drift and, moreover, constrained to lie within the conic region (blue dotted lines) according to (7). Notice that as the starting point of the geodesics approaches the singularity at $r=0$, the displacement of the center of the unit sphere increases, resulting in more acute cones. In the vicinity of the central singularity, all geodesics are rapidly dragged towards it, and there is no way to prevent them from reaching $r=0$, in perfect agreement with the understanding that $r=0$ is a future timelike singularity.
@@ III.1 Static black holes
For $r=r_{+}$ and $r=r_{-}$, we have $f(r)=0$, which correspond to the critical Zermelo case. The associated geodesics are such that $2\leq dx\leq 0$, as in the case of the Schwarzschild event horizon. In the region between the two Killing horizons, $r_{-}<r<r_{+}$, where $f(r)<0$, we have the strong drift case. This regime is similar to the one occurring in the interior region of the Schwarzschild black hole, where we have only ingoing geodesics since $dx<0$. The novelty for the Reissner-Nordström black hole is that, for $r_{0}\leq r<r_{-}$, we have again the weak drift regime corresponding to $f(r)>0$. In this regime, both ingoing ($dx<0$) and outgoing ($dx>0$) null geodesics are allowed, but they remain entirely confined within the inner horizon. This indicates that null geodesics starting within the inner (Cauchy) horizon of the Reissner-Nordström black hole can escape from reaching the spacelike singularity at $r=0$. This characteristic is well-known in the theoretical study of charged black holes. However, it is generally not considered physically relevant because the electric charge in astrophysical black holes is expected to neutralize quickly [34, 35]. Additionally, there are inherent instabilities associated with the Cauchy horizon, see [36, 37, 38].
@@
We remark that the Painlevé-Gullstrand coordinates for the Reissner-Nordström black hole are defined only for $r>r_{0}$. Consequently, the Finslerian description of its causal structure is limited by the same restriction. However, this does not pose a major concern for our purposes, as it is related to the presence of a negative local mass, which effectively prevents ingoing geodesics from reaching the central singularity at $r=0$, a well-known issue with the Painlevé-Gullstrand coordinates. For further details, see [39]. For an interpretation of this phenomenon in the context of the river model of black holes, refer to [25].

### v2

# Wind-Finslerian structure of black holes

@@ Abstract
Recently, there has been an increasing interest in the Finslerian interpretation of null geodesics in the exterior regions of stationary black holes, particularly through the Zermelo navigation problem and the Randers metric. In this work, we show that recent mathematical advancements in wind-Finslerian structures, which involve the critical and strong Zermelo navigation problems and their connections to Kropina and Lorentz-Finsler metrics, enable the extension of the Finslerian framework to encompass horizons and their interior regions of black holes. The Finslerian indicatrix, a key element of this framework, serves as an effective tool for identifying frame-dragging effects and the location of horizons and ergosurfaces. We illustrate our results with explicit physical examples, focusing on spherically symmetric black holes, Kerr black holes, and analog models of gravity. Our findings provide new insights into the “river model” of black holes, offering enhanced visual representations of null geodesics on the ergosurfaces, horizons, and within their interior regions.
@@ I Introduction
where $\mathbf{x}$ is a point with coordinates $(x^{\sigma})$ in the spacetime manifold $\mathcal{M}$ and $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$. We denote the components of $d\mathbf{x}$ by $(dx^{\mu})$, following the standard physical convention in which Greek indices run over all $n+1$ spacetime coordinates, while Latin indices are limited to the $n$ spatial dimensions. In the Randers metric (1), $a_{\mu\nu}$ is a standard metric tensor (Lorentzian, in the case of Spacetime Physics), and $b_{\mu}$ is a one-form (covariant vector) that encodes a directional dependence of the geometry. If $b_{\mu}=0$, the underlying geometry reduces to the standard Riemannian/Lorentzian one associated with the metric tensor $a_{\mu\nu}$. Notice that $F(\mathbf{x},d\mathbf{x})=F(\mathbf{x},-d\mathbf{x})$ for generic vectors $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$ if, and only if, $b_{\mu}=0$. In other words, a Randers metric is symmetric if and only if it is Riemannian. Notwithstanding, generic Finsler metrics are positively homogeneous of degree one, i.e., $F(\mathbf{x},\lambda d\mathbf{x})=\lambda F(\mathbf{x},d\mathbf{x})$ for all $\lambda\geq 0$. All Finslerian-geometric objects derived from (1), such as the fundamental form, connections, covariant derivative, and curvature, are defined starting from $F^{2}(\mathbf{x},d\mathbf{x})$ analogously to the Riemannian case (see [2] and references therein for further details). Positive definite metrics of the Randers type (1) arise naturally [5] in the classical optimization problem known as the Zermelo navigation problem [6], which consists in finding minimal (non-relativistic) time trajectories in curved spaces under the influence of vector drifts (winds). Many recent advancements in Finsler geometry and its applications in Physics are closely related to the Zermelo problem and its solutions involving metrics such as those of Randers type. The central mathematical concept explored in the present paper is the notion of a wind-Finslerian structure, which provides a unified model for the classical Zermelo navigation problem across different types of drifts. This concept was introduced and developed in [7]. To be more precise, in all particular cases considered here, the indicatrices are ellipsoids and, consequently, we are dealing with the specific class of wind-Riemannian structures, according to the definitions of [7]. In this paper, we will further explore the wind-Finslerian structures and their connections to concrete examples in Spacetime Physics. The reference [7] also provides a comprehensive review of the recent mathematical literature on this topic.
@@
From a physical standpoint, one does not expect distinguished local spacelike directions. However, this does not apply to timelike directions; in many circumstances, the future and the past causal cones are not merely related by simple inversions. In fact, this was the main motivation behind Randers’ introduction of a spacetime metric with asymmetrical properties of the type (1) in the early forties [4]. Moreover, the Principle of Equivalence imposes restrictions on possible forms of a Finsler metric from the Spacetime Physics perspective [8]. Interestingly, positive definite metrics of type (1) often appear in specific spacetime problems restricted to spatial sections, such as the description of causal structures. In particular, the propagation of light rays in stationary spacetimes can be effectively described by a Randers spatial metric. These issues were mathematically established in [9] and then considered from a physical point of view for black holes in [10]. The extension for nonstationary but stably causal spacetimes admitting a globally defined time coordinate was considered in [11]. More recently, [12] has shown that the Zermelo problem can be used to generate spacetimes with specific causal properties. To date, only the propagation of light rays in the exterior region of black holes with a non-vanishing Killing vector has been described in Finslerian terms in the Physics literature. Extending the Finsler structure across the event horizon is a physically relevant problem, as recognized in [12]. Notably, recent research in the Mathematics literature has tackled these issues. For instance, [13] shows that there exists a one-to-one correspondence between Randers spaces and conformal classes of stationary spacetimes, with the causal structure of the latter fully encoded in the metric properties of the former. Furthermore, the primary objective of Ref. [7] was to extend the Finslerian description of causality and horizons in any region of a spacetime admitting a spacelike slice and a non-vanishing transversal Killing vector field. One of the main goals of the current study is precisely to establish a connection between these recent mathematical findings and the physical literature. We highlight also that recent applications of Finsler geometry in other problems of General and Special Relativity can be found, e.g., in [14, 15, 16, 17, 18, 19, 20, 21, 22].
@@
In this paper, we demonstrate how recent mathematical progress [7, 9, 13, 23, 24, 25] can be explored to interpret light propagation within the interior region of black holes in Finslerian terms. The key ingredients are the Kropina and Lorentz-Finsler metrics, which appear in certain extended versions of the Zermelo navigation problem. We show, through explicit and physically relevant examples, that the Kropina metric adequately describes the causal properties of ergosurfaces in stationary black holes, while the Lorentz-Finsler metric captures, in the region where it is well defined, the causal structure of the interior (non-stationary) region of black holes. In this way, we show that it is indeed possible to have, based on recent mathematical developments, a complete Finslerian description of the causal structure of black holes, shedding new light on phenomena such as the one-directional membrane behavior of horizons [26] and the river model paradigm of black holes [27].
@@
The paper is organized as follows. In the next section, we review recent mathematical developments related to the Kropina and Lorentz-Finsler metrics and compile the results that are instrumental in describing the causal structure of black holes, which is explored in Section III. In Sec. IV, we analyze analog models of gravity within the Finslerian framework. The final section is dedicated to our concluding remarks, which include brief comments on Fermat’s principle and analog spacetimes. We adopt $G=c=1$ units throughout this work.
@@ II Finsler geometry and the Zermelo problem
where $a^{ij}$ is the inverse of the metric tensor $a_{ij}$. Hence, the weak drift also implies $||b||^{2}_{F}<1$, which is commonly known as the Finsler condition in the Physics literature. The so-called Randers $(a_{ij},b_{j})$ and Zermelo $(h_{ij},W^{j})$ data are connected through a Legendre transformation [10]. As it will become clear in the next section, describing the wind-Finslerian structure of black holes requires extending the Zermelo problem to the critical $(||W||=1)$ and strong $(||W||>1)$ drift cases. Fortunately, both problems have been thoroughly investigated and, for the purposes of this work, conclusively solved in the recent mathematical literature. Ref. [7], which first solved the strong drift case, also provides a useful review of the pertinent mathematical literature.
@@
The strong drift case, on the other hand, was solved in [7], where it was shown that minimal time trajectories are associated with the geodesics of the minus-sign case of the so-called Lorentz-Finsler metrics
@@
We shall refer to the minus and plus cases, respectively, as first and second Lorentz-Finsler metrics. The strong drift condition $\lambda<0$ guarantees that both metrics in (6) are positive functions on $A$. Additionally, as in the critical case, the problem exhibits a one-directional behavior: a solution exists only if the trajectories are positively aligned with the drift, and moreover their instantaneous velocities lie within the conic region (7). It is worth noticing that the transition from weak to strong drift involves several mathematical subtleties. We will adopt a more heuristic approach in the next subsection. The reader can find the pertinent mathematical details in Ref. [7].
@@
In other words, the indicatrices of the Finsler metrics associated with the Zermelo navigation problem on a Riemannian manifold $(\mathcal{M},h_{ij})$ with drift $W^{i}$, regardless of the value of $||W||$, correspond to the unit balls of the Riemannian metric $h_{ij}$, with their centers shifted by the vector $W^{i}$. Randers had already recognized this type of eccentricity in his pioneering work [4]. The use of displaced indicatrices as tools to describe causal and geometric structure, an important result of Ref. [7], will be essential for the analysis of the causal structure of black holes in Section III.
@@ II.1 Geodesic matching
The mathematical description of the causal structure of Finsler spaces is both comprehensive and sophisticated, see, e.g.[7, 28, 29, 30, 31]. For our purposes here, we need only to consider the matching of geodesics across the three cases of Zermelo’s problems, whose general validity is assured by the rigorous mathematical results on wind-Finslerian structures introduced in [7]. Differentiability issues will not be discussed, as all functions are assumed to be sufficiently smooth. Let $W^{i}$ be a generic smooth vector field drift on a Riemannian manifold $(\mathcal{M},h_{ij})$ and assume there are regions of $\mathcal{M}$ characterized by weak ($||W||<1$) and strong $(||W||>1)$ drifts. The boundary between these regions corresponds to the critical drift case $||W||=1$, and in the physical cases where it is associated with a horizon, it will be a co-dimension 1 submanifold $\mathcal{H}$ of $\mathcal{M}$. The key observation is that geodesics can cross smoothly from the weak to the strong drift regions, provided that the half-space (5) and conic (7) constraints are satisfied in the pertinent regions. This is possible because the Randers (1), Kropina (4), and first Lorentz-Finsler (6) metrics all give rise to the same Finslerian Lagrangian action $S=\int F(\mathbf{x},d\mathbf{x})$ for geodesics, and they can be smoothly matched across the drift region transitions. In fact, the three cases can be nicely unified if one uses the Finslerian Lagrangian action with
$F(\mathbf{x},d\mathbf{x})=\frac{h_{ij}dx^{i}dx^{j}}{W_{i}dx^{i}+\sqrt{\left(W_{j}dx^{j}\right)^{2}+\lambda h_{ij}dx^{i}dx^{j}}},$ (9)
which is valid for any value of $\lambda$, see [7] for further details. Notice that the half-space and conic constraints are essential to assure that we have positive Lagrangians $F(\mathbf{x},d\mathbf{x})$ and consequently well-defined geodesic problems, on $\mathcal{H}$ and in the strong drift region.
@@
Figure 1: Depiction of the Finslerian indicatrices for the three drift regimes of the Schwarzschild case along the $x$ axis in the equatorial plane. In all cases, the indicatrices are unit spheres, with their centers displaced in the $x$ direction according to (11). Right: The weak drift regime $r>2M$, in which null geodesics starting from a given point can propagate in all directions. Center: The critical regime $r=2M$, in which the propagation of null geodesics is restricted to the directions that are positively aligned with the drift according to the half-space restriction (5). In this case, the half-space is tangent to the critical submanifold $\mathcal{H}$ (the event horizon), implying that there is no outgoing null geodesics starting on $\mathcal{H}$, establishing its causal disconnection from the exterior region. Left: The strong drift regime. Here, propagation is similarly limited to directions positively aligned with the drift and, moreover, constrained to lie within the conic region (blue dotted lines) according to (7). Notice that as the starting point of the geodesics approaches the singularity at $r=0$, the displacement of the center of the unit sphere increases, resulting in more acute cones. In the vicinity of the central singularity, all geodesics are rapidly dragged towards it, and there is no way to prevent them from reaching $r=0$, in perfect agreement with the understanding that $r=0$ is a future timelike singularity. This is a concrete realization, for a Schwarzschild black hole, of the wind-Finslerian description of the causal structure of spacetimes introduced in [7].
@@ III.1 Static black holes
For $r=r_{+}$ and $r=r_{-}$, we have $f(r)=0$, which correspond to the critical Zermelo case. The associated geodesics are such that $-2\leq dx\leq 0$, as in the case of the Schwarzschild event horizon. In the region between the two Killing horizons, $r_{-}<r<r_{+}$, where $f(r)<0$, we have the strong drift case. This regime is similar to the one occurring in the interior region of the Schwarzschild black hole, where we have only ingoing geodesics since $dx<0$. The novelty for the Reissner-Nordström black hole is that, for $r_{0}\leq r<r_{-}$, we have again the weak drift regime corresponding to $f(r)>0$. In this regime, both ingoing ($dx<0$) and outgoing ($dx>0$) null geodesics are allowed, but they remain entirely confined within the inner horizon. This indicates that null geodesics starting within the inner (Cauchy) horizon of the Reissner-Nordström black hole can escape from reaching the spacelike singularity at $r=0$. This characteristic is well-known in the theoretical study of charged black holes. However, it is generally not considered physically relevant because the electric charge in astrophysical black holes is expected to neutralize quickly [35, 36]. Additionally, there are inherent instabilities associated with the Cauchy horizon, see [37, 38, 39].
@@
We remark that the Painlevé-Gullstrand coordinates for the Reissner-Nordström black hole are defined only for $r>r_{0}$. Consequently, the Finslerian description of its causal structure is limited by the same restriction. However, this does not pose a major concern for our purposes, as it is related to the presence of a negative local mass, which effectively prevents ingoing geodesics from reaching the central singularity at $r=0$, a well-known issue with the Painlevé-Gullstrand coordinates. For further details, see [40]. For an interpretation of this phenomenon in the context of the river model of black holes, refer to [27]. Notwithstanding, Finsler geometry offers a fresh and physically insightful view of this issue. The local mass function relevant to the problem is the so-called Misner-Sharp-Hernandez mass $M_{\rm MSH}(r)$, which plays a central role in spherical fluid mechanics and gravitational collapse [41, 42]. For a static spherically symmetrical spacetime, this mass is introduced by recasting the function $f$ in (14) as
$f(r)=1-\frac{2M_{\rm MSH}(r)}{r},$ (22)
this mimicking some properties of the Schwarzschild mass. For a Reissner-Nordström black hole, this local mass reads
$M_{\rm MSH}(r)=M-\frac{Q^{2}}{2r},$ (23)
implying that $M_{\rm MSH}$ is non-positive for $0<r\leq r_{0}$, thus resulting in a non-attractive behavior that prevents all geodesics from reaching the center $r=0$. Remarkably, when expressed in terms of the Misner–Sharp–Hernandez mass, the Finsler metric (18) for generic static, spherically symmetric spacetimes with $R(r)=r$, along with the associated indicatrix expressions (19) and (20), reduce to those of the Schwarzschild case (10), (11), and (12), with the Schwarzschild mass replaced by $M_{\rm MSH}(r)$. This shows that $M_{\rm MSH}(r)$ does more than mimic the Schwarzschild mass in the underlying Finsler geometry of a static spherically symmetric black hole; it fully encodes all the Finsler geometrical properties, offering a new geometric perspective in the Misner-Sharp-Hernandez mass.
Yet, regarding the issue of geodesic repulsion in the region $0<r\leq r_{0}$, one should recall the Martel-Poisson family of coordinates [32, 40], which can effectively describe free-fall motion for $r<r_{0}$ and could thus be explored to extend the wind-Finslerian structure of black holes to regions with negative Misner-Sharp-Hernandez mass. Such coordinates correspond to the choice
$dt=d\tau+\frac{1}{f}\sqrt{1-pf}dr,$ (24)
instead of (15), with constant $p$ such that $0<p\leq 1$. For the physical interpretation of the parameter $p$, see [32, 40]. Notice that $p=1$ corresponds to the usual Painlevé-Gullstrand coordinates and the (singular) limit $p\to 0$ is related to the Eddington-Finkelstein coordinates [40]. Inserting (24) in (14) and looking for forward-in-time null geodesics, one obtains the Finsler metric
$\!dt\!=\!F(x,dx)\!=\!\frac{1}{f}\left(\!\sqrt{1-pf}dr+\sqrt{dr^{2}+fR^{2}dS^{2}}\right)\!,\!$ (25)
which, in the Reissner-Nordström case, is associated to the ellipsoidal indicatrix
$p\left(dx+\frac{\sqrt{1-pf}}{p}\right)^{2}+dy^{2}+dz^{2}=\frac{1}{p},$ (26)
implying
$-1-\sqrt{1-pf}\leq p\,dx\leq 1-\sqrt{1-pf}.$ (27)
From (3) and (25), we see that the critical drift region is unaltered by $p$; it always corresponds to the zeros of $f(r)$, and hence the qualitative behavior of the null geodesics obeying (26) and (27) remains the same regardless of the value of the constant $p$. However, the Finsler structure associated with the Martel-Poisson family of coordinates is well defined for any $r>r_{0}^{*}$, with $r_{0}^{*}$ being the solution of
$\frac{2M_{\rm MSH}(r)}{r}=-\frac{1-p}{p}.$ (28)
Note that $r_{0}^{*}$ lies within the region $0<r\leq r_{0}$ of geodesic repulsion, where $M_{\rm MSH}$ is negative. Finsler geometry, consequently, plays a significant role in understanding the internal region of static black holes, enlightening the relationship between the Misner-Sharp-Hernandez local mass function and the causal structure of the spacetime.